# Metric3D v2: A Versatile Monocular Geometric Foundation Model for Zero-shot Metric Depth and Surface Normal Estimation

### v1

@@ Abstract
We introduce Metric3D v2, a geometric foundation model for zero-shot metric depth and surface normal estimation from a single image, which is crucial for metric 3D recovery. While depth and normal are geometrically related and highly complimentary, they present distinct challenges. State-of-the-art (SoTA) monocular depth methods achieve zero-shot generalization by learning affine-invariant depths, which cannot recover real-world metrics. Meanwhile, SoTA normal estimation methods have limited zero-shot performance due to the lack of large-scale labeled data. To tackle these issues, we propose solutions for both metric depth estimation and surface normal estimation. For metric depth estimation, we show that the key to a zero-shot single-view model lies in resolving the metric ambiguity from various camera models and large-scale data training. We propose a canonical camera space transformation module, which explicitly addresses the ambiguity problem and can be effortlessly plugged into existing monocular models. For surface normal estimation, we propose a joint depth-normal optimization module to distill diverse data knowledge from metric depth, enabling normal estimators to learn beyond normal labels. Equipped with these modules, our depth-normal models can be stably trained with over $16$ million of images from thousands of camera models with different-type annotations, resulting in zero-shot generalization to in-the-wild images with unseen camera settings. Our method current [^1] Feb. tbd,2024 ranks the 1st on various zero-shot and non-zero-shot benchmarks for metric depth, affine-invariant-depth as well as surface-normal prediction, shown in Fig. 1. Notably, we surpassed the ultra-recent MarigoldDepth and DepthAnything on various depth benchmarks including NYUv2 and KITTI. Our method enables the accurate recovery of metric 3D structures on randomly collected internet images, paving the way for plausible single-image metrology. The potential benefits extend to downstream tasks, which can be significantly improved by simply plugging in our model. For example, our model relieves the scale drift issues of monocular-SLAM (Fig. 3), leading to high-quality metric scale dense mapping. These applications highlight the versatility of Metric3D v2 models as geometric foundation models.
Our project page is at https://JUGGHM.github.io/Metric3Dv2.
@@ 1 Introduction
With the proposed method, we can stably scale up model training to $16$ million images from 18 datasets of diverse scene types (indoor and outdoor, real or synthetic data), camera models (tens of thousands of different cameras), and annotation categories (with or without normal), leading to zero-shot transferability and significantly improved accuracy. Fig. 4 illustrates how the large-scale data with depth annotations directly facilitate metric depth and surface normal learning. The metric depth and normal given by our model directly broaden the applications in downstream tasks. We achieve state-of-the-art performance on over 16 depth and normal benchmarks, see Fig. 1. Our model can accurately reconstruct metric 3D from randomly collected Internet images, enabling plausible single-image metrology. As an example (Fig. 3), with the predicted metric depths from our model, we significantly reduce the scale drift of monocular SLAM [37, 47] systems, achieving much better mapping quality with real-world metric recovery. Our model also enables large-scale 3D reconstruction [48]. To summarize, our main contributions are:
  • We propose a canonical and de-canonical camera transformation method to solve the metric depth ambiguity problems from various camera settings. It enables the learning of strong zero-shot monocular metric depth models from large-scale datasets.
  • We propose a random proposal normalization loss to effectively boost the depth accuracy;
@@ 2 Related Work
3D reconstruction from a single image. Reconstructing various objects from a single image has been well studied [51, 52, 53]. They can produce high-quality 3D models of cars, planes, tables, and human body [54, 55]. The main challenge is how to best recover objects’ details, how to represent them with limited memory, and how to generalize to more diverse objects. However, all these methods rely on learning priors specific to a certain object class or instance, typically from 3D supervision, and can therefore not work for full scene reconstruction. Apart from these reconstructing objects works, several works focus on scene reconstruction from a single image. Saxena et al. [56] construct the scene based on the assumption that the whole scene can be segmented into several small planes. With planes’ orientation and location, the 3D structure can be represented. Recently, LeReS [25] proposed to use a strong monocular depth estimation model to do scene reconstruction. However, they can only recover the shape up to a scale. Zhang et al. [57] recently proposed a zero-shot geometry-preserving depth estimation model that is capable of making depth predictions up to an unknown scale, without requiring scale-invariant depth annotations for training. In contrast to these works, our method can recover the metric 3D structure.
@@
Supervised monocular depth estimation. After several benchmarks [58, 59] are established, neural network based methods [17, 30, 19] have dominated since then. Several approaches regress the continuous depth from the aggregation of information in an image [60]. As depth distribution corresponding to different RGBs can vary to a large extent, some methods [30, 19] discretize the depth and formulate this problem to a classification [18], which often achieves better performance. The generalization issue of deep models for 3D metric recovery is related to two problems. The first one is to generalize to diverse scenes, while the other one is how to predict accurate metric information under various camera settings. The first problem has been well addressed by recent methods. Some works [22, 21, 18] propose to construct a large-scale relative depth dataset, such as DIW [24] and OASIS [23], and then they target learning the relative relations. However, the relative depth loses geometric structure information. To improve the recovered geometry quality, learning affine-invariant depth methods, such as MiDaS [27], LeReS [25], and HDN [29] are proposed. By mixing large-scale data, state-of-the-art performance and the generalization over scenes are improved continuously. Note that by design, these methods are unable to recover the metric information. How to achieve both strong generalization and accurate metric information over diverse scenes is the key problem that we attempt to tackle.
@@
Surface normal estimation. Compared to metric depth, surface normal suffers no metric ambiguity and preserves local geometry better. These properties attract researchers to apply normal in various vision tasks like localization [11], mapping [61], and 3D scene reconstruction [6, 62]. Currently, learning-based methods [33, 34, 42, 63, 64, 65, 66, 32, 67, 68, 69, 70, 62] have dominated monocular surface normal estimation. Since normal labels required for training cannot be directly captured by sensors, previous works use [58, 65, 41, 63] kernel functions to annotate normal from dense indoor depth maps [58]. These annotations become incomplete on reflective surfaces and inaccurate at object boundaries. To learn from such imperfect annotations, GeoNet [41] proposes to enforce depth-normal consistency with mutual transformation modules, ASN [70, 69] propose a novel adaptive surface normal constraint to facilitate joint depth-normal learning, and Bae et al. [33] propose an uncertainty-based learning objective. Nonetheless, it is challenging for such methods to further increase their generalization, due to the limited dataset size and the diversity of scenes, especially for outdoor scenarios. Omni-data [35] advances to fill this gap by building $1300$M frames of normal annotation. Normal-in-the-wild [71] proposes a pipeline for efficient normal labeling. However, further scaling up normal labels remains difficult. This underscores research significance in finding an efficient way to distill prior from other types of annotation.
@@
Deep iterative refinement for geometry. Iterative refinement enables multi-step coarse-to-fine prediction and benefits a wide range of geometry estimation tasks, such as optical flow estimation [72, 46, 73], depth completion [74, 43, 75], and stereo matching [45, 76, 77]. Classical iterative refinements [72, 74] optimize directly on high-resolution outputs using high-computing-cost operators, limiting researchers from applying more iterations for better predictions. To address this limitation, RAFT[46] proposes to optimize an intermediate low-resolution prediction using ConvGRU modules. For monocular depth estimation, IEBins[44] employs similar methods to optimize depth-bin distribution. Differently, IronDepth [42] propagates depth on pre-computed local surfaces. Regarding surface normal refinement, Lenssen et al. [78] propose a deep iterative method to optimize normal from point clouds. Zhao et al. [79] design a solver to refine depth and normal jointly, but it requires multi-view prior and per-sample post optimization. Without multi-view prior, such a non-learnable optimization method could fail due to unsatisfactory initial predictions. All the monocular methods [44, 42, 78], however, iterate over either depth or normal independently. In contrast, our joint optimization module tightly couples depth and normal with each other.
@@
Large-scale data training. Recently, various natural language problems and computer vision problems [80, 81, 82] have achieved impressive progress with large-scale data training. CLIP [81] is a promising classification model, which is trained on billions of paired image and language description data. It achieves state-of-the-art performance over several classification benchmarks by zero-shot testing. Dinov2 [83] collects $142$M images to conduct vision-only self-supervised learning for vision transformers[84]. Generative models like LDM [85] have also undergone billion-level data pre-training. For depth prediction, large-scale data training has been widely applied. Ranft et al. [27] mix over 2 million data in training, LeReS [26] collects over $300$ thousands data, Eftekhar et al. [35] also merge millions of data to build a strong depth prediction model. For surface normal estimation, Omni-data [35] performs dense reconstruction to generate $14$M frames with surface normal annotations, aggregating data from six different datasets. Furthermore, it established multi-task training involving these annotations.
@@ 3.1 Ambiguity Issues in Metric Depth Estimation
Fig. 5 presents an example of photos taken by different cameras and at different distances. Only from the image’s appearance, one may think the last two photos are taken at a similar location by the same camera. In fact, due to different focal lengths, these are captured at different locations. Thus, camera intrinsic parameters are critically important for the metric estimation from a single image, as otherwise, the problem is ill posed. To avoid such metric ambiguity, recent methods, such as MiDaS [27] and LeReS [25], decouple the metric from the supervision and compromise learning the affine-invariant depth.
@@
Fig. 6 (A) shows a simple pin-hole perspective projection. Object $A$ locating at $d_{a}$ is projected to $A^{\prime}$. Based on the principle of similarity, we have the equation:
@@
where $\hat{S}$ and $\hat{S^{\prime}}$ are the real and imaging size respectively. $\hat{\cdot}$ denotes variables are in the physical metric (e.g., millimeter). To recover $d_{a}$ from a single image, focal length, imaging size of the object, and real-world object size must be available. Estimating the focal length from a single image is a challenging and ill-posed problem. Although several methods [25, 86] have explored, the accuracy is still far from being satisfactory. Here, we simplify the problem by assuming the focal length of a training/test image is available. In contrast, understanding the imaging size is much easier for a neural network. To obtain the real-world object size, a neural network needs to understand the semantic scene layout and the object, at which a neural network excels. We define $\alpha=\nicefrac{{\hat{f}}}{{\hat{S^{\prime}}}}$, so $d_{a}$ is proportional to $\alpha$.
Figure 6: Pinhole camera model. (A) Object $A$ at the distance $d_{a}$ is projected to the image plane. (B) Using two cameras to capture the car. The left one has a larger pixel size. Although the projected imaging sizes are the same, the pixel-represented images (resolution) are different.
We make the following observations regarding sensor size, pixel size, and focal length.
O1: Sensor size and pixel size do not affect the metric depth estimation. Based on the perspective projection (Fig. 6 (A)), the sensor size only affects the field of view (FOV) and is irrelevant to $\alpha$, thus does not affect the metric depth estimation. For the pixel size, we assume two cameras with different pixel sizes ($\delta_{1}=2\delta_{2}$) but the same focal length $\hat{f}$ to capture the same object locating at $d_{a}$. Fig. 6 (B) shows their captured photos. According to the preliminaries, the pixel-represented focal length $f_{1}=\frac{1}{2}f_{2}$. As the second camera has a smaller pixel size, although in the same projected imaging size $\hat{S^{\prime}}$, the pixel-represented image resolution is $S^{\prime}_{1}=\frac{1}{2}S^{\prime}_{2}$. According to Eq. (1), $\frac{\hat{f}}{\delta_{1}\cdot S^{\prime}_{1}}=\frac{\hat{f}}{\delta_{2}\cdot S^{\prime}_{2}}$, i.e. $\alpha_{1}=\alpha_{2}$, so $d_{1}=d_{2}$. Therefore, different camera sensors would not affect the metric depth estimation.
O2: The focal length is vital for metric depth estimation. Fig. 5 shows the metric ambiguity issue caused by the unknown focal length. Fig. 7 illustrates this. If two cameras ($\hat{f}_{1}=2\hat{f}_{2}$) are at distances $d_{1}=2d_{2}$, the imaging sizes on cameras are the same. Thus, only from the appearance, the network will be confused when supervised with different labels. Based on this observation, we propose a canonical camera transformation method to solve the supervision and image appearance conflicts.
@@
Unlike depth, surface normal does not have any metric ambiguity problem. In Fig. 8, we illustrate this concept with two depth maps at varying scales, denoted as $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, featuring distinct metrics d1 and d2, respectively, where $d_{1}$ $d_{2}$. After upprojecting the depth to the 3D point cloud, the dolls are in different distances, i.e. $d_{1}$ and $d_{2}$ and $d_{1}<d_{2}$. However, despite the variation in depth scales, the surface normals $\mathbf{n}_{1}$ and $\mathbf{n}_{2}$ corresponding to a certain pixel $A^{\prime}\in\mathbf{I}$ remain the same.

### v4

@@ Abstract
We introduce Metric3D v2, a geometric foundation model for zero-shot metric depth and surface normal estimation from a single image, which is crucial for metric 3D recovery. While depth and normal are geometrically related and highly complimentary, they present distinct challenges. State-of-the-art (SoTA) monocular depth methods achieve zero-shot generalization by learning affine-invariant depths, which cannot recover real-world metrics. Meanwhile, SoTA normal estimation methods have limited zero-shot performance due to the lack of large-scale labeled data. To tackle these issues, we propose solutions for both metric depth estimation and surface normal estimation. For metric depth estimation, we show that the key to a zero-shot single-view model lies in resolving the metric ambiguity from various camera models and large-scale data training. We propose a canonical camera space transformation module, which explicitly addresses the ambiguity problem and can be effortlessly plugged into existing monocular models. For surface normal estimation, we propose a joint depth-normal optimization module to distill diverse data knowledge from metric depth, enabling normal estimators to learn beyond normal labels. Equipped with these modules, our depth-normal models can be stably trained with over $16$ million of images from thousands of camera models with different-type annotations, resulting in zero-shot generalization to in-the-wild images with unseen camera settings. Our method currently [^1] Feb. tbd,2024 ranks the 1st on various zero-shot and non-zero-shot benchmarks for metric depth, affine-invariant-depth as well as surface-normal prediction, shown in Fig. 1. Notably, we surpassed the ultra-recent MarigoldDepth and DepthAnything on various depth benchmarks including NYUv2 and KITTI. Our method enables the accurate recovery of metric 3D structures on randomly collected internet images, paving the way for plausible single-image metrology. The potential benefits extend to downstream tasks, which can be significantly improved by simply plugging in our model. For example, our model relieves the scale drift issues of monocular-SLAM (Fig. 3), leading to high-quality metric scale dense mapping. These applications highlight the versatility of Metric3D v2 models as geometric foundation models. Our project page is at https://JUGGHM.github.io/Metric3Dv2.
@@ 1 Introduction
With the proposed method, we can stably scale up model training to $16$ million images from 16 datasets of diverse scene types (indoor and outdoor, real or synthetic data), camera models (tens of thousands of different cameras), and annotation categories (with or without normal), leading to zero-shot transferability and significantly improved accuracy. Fig. 4 illustrates how the large-scale data with depth annotations directly facilitate metric depth and surface normal learning. The metric depth and normal given by our model directly broaden the applications in downstream tasks. We achieve state-of-the-art performance on over 16 depth and normal benchmarks, see Fig. 1. Our model can accurately reconstruct metric 3D from randomly collected Internet images, enabling plausible single-image metrology. For examples (Fig. 3), we recovery real-world metric to improve monocular SLAM [37, 47] and facilitate large-scale 3D reconstruction [48]. Our main contributions can be summarized as:
  • We propose a canonical camera transformation method to address metric depth ambiguity across different camera settings. This approach facilitates training zero-shot monocular metric depth models using large-scale datasets.
  • We design a random proposal normalization loss to effectively improve metric depth.
@@ 2 Related Work
3D reconstruction from a single image. The reconstruction of diverse objects from a singular image has been extensively investigated in prior research [51, 52, 53]. These methodologies exhibit proficiency in generating high-fidelity 3D models encompassing various entities such as cars, planes, tables, and human bodies [54, 55]. The primary challenge lies in optimizing the recovery of object details, devising efficient representations within constrained memory resources, and achieving generalization across a broader spectrum of objects. However, these approaches typically hinge upon learning object-specific or instance-specific priors, often derived from 3D supervision, thereby rendering them unsuitable for comprehensive scene reconstruction. In addition to the aforementioned efforts on object reconstruction, several studies focus on scene reconstruction from single images. Saxena et al.[56] adopt an approach that segments the entire scene into multiple small planes, with the 3D structure represented based on the orientation and positioning of these planes. More recently, LeReS[25] proposed employing a robust monocular depth estimation model for scene reconstruction. Nonetheless, their method is limited to recovering shapes up to a certain scale. Zhang et al. [57] recently introduced a zero-shot geometry-preserving depth estimation model capable of providing depth predictions up to an unknown scale. In contrast to the aforementioned methodologies, our approach excels in recovering the metric 3D structure of scenes.
@@
Supervised monocular depth estimation. Following the establishment of several benchmarks [58, 59], neural network-based methods [17, 30, 19] have dominated this task. These approaches often regress continuous depth by aggregating the information from an image [60]. However, since depth distribution varies significantly with different RGB values, some methods tend to discretize depth and reformulate the problem as a classification task [18] for better performance. The generalization of deep models for 3D metric recovery faces two challenges: adapting to diverse scenes and predicting accurate metric information under various camera settings. Recent methods [22, 21, 18] have effectively addressed the first challenge by creating large-scale relative depth datasets like DIW [24] and OASIS [23] to learn relative relations, which lose geometric structure information. To enhance geometry, methods like MiDaS [27], LeReS [25], and HDN [29] employ affine-invariant depth learning. These approaches, utilizing large-scale data, have continuously improved performance and scene generalization. However, they inherently struggle to recover metric information. Thus, achieving both strong generalization and accurate metric data across diverse scenes remains a key challenge to be addressed. Con-currently, ZoeDepth [36], ZeroDepth [61], and UniDepth [62] apply varying strategies to tackle this challenge.
@@
Surface normal estimation. Compared to metric depth, surface normal suffers no metric ambiguity and preserves local geometry better. These properties attract researchers to apply normal in various vision tasks like localization [11], mapping [63], and 3D scene reconstruction [6, 64]. Currently, learning-based methods [33, 34, 42, 65, 66, 67, 68, 32, 69, 70, 71, 72, 64] have dominated monocular surface normal estimation. Since normal labels required for training cannot be directly captured by sensors, previous works use [58, 67, 41, 65] kernel functions to annotate normal from dense indoor depth maps [58]. These annotations become incomplete on reflective surfaces and inaccurate at object boundaries. To learn from such imperfect annotations, GeoNet [41] proposes to enforce depth-normal consistency with mutual transformation modules, ASN [72, 71] propose a novel adaptive surface normal constraint to facilitate joint depth-normal learning, and Bae et al. [33] propose an uncertainty-based learning objective. Nonetheless, it is challenging for such methods to further increase their generalization, due to the limited dataset size and the diversity of scenes, especially for outdoor scenarios. Omni-data [35] advances to fill this gap by building $1300$M frames of normal annotation. Normal-in-the-wild [73] proposes a pipeline for efficient normal labeling. A con-current work DSINE [74] also employs diverse datasets to train generalizable surface normal estimators. However, further scaling up normal labels remains difficult. This underscores research significance in finding an efficient way to distill prior from other types of annotation.
@@
Deep iterative refinement for geometry. Iterative refinement enables multi-step coarse-to-fine prediction and benefits a wide range of geometry estimation tasks, such as optical flow estimation [75, 46, 76], depth completion [77, 43, 78], and stereo matching [45, ma2022multiview, 79]. Classical iterative refinements [75, 77] optimize directly on high-resolution outputs using high-computing-cost operators, limiting researchers from applying more iterations for better predictions. To address this limitation, RAFT[46] proposes to optimize an intermediate low-resolution prediction using ConvGRU modules. For monocular depth estimation, IEBins[44] employs similar methods to optimize depth-bin distribution. Differently, IronDepth [42] propagates depth on pre-computed local surfaces. Regarding surface normal refinement, Lenssen et al. [80] propose a deep iterative method to optimize normal from point clouds. Zhao et al. [81] design a solver to refine depth and normal jointly, but it requires multi-view prior and per-sample post optimization. Without multi-view prior, such a non-learnable optimization method could fail due to unsatisfactory initial predictions. All the monocular methods [44, 42, 80], however, iterate over either depth or normal independently. In contrast, our joint optimization module tightly couples depth and normal with each other.
@@
Large-scale data training. Recently, various natural language problems and computer vision problems [82, 83, 84] have achieved impressive progress with large-scale data training. CLIP [83] is a promising classification model trained on billions of paired image-language data pairs, achieving achieves state-of-the-art performance zero-shot classification benchmarks. Dinov2 [85] collects $142$M images to conduct vision-only self-supervised learning for vision transformers[86]. Generative models like LDM [87] have also undergone billion-level data pre-training. For depth prediction, large-scale data training has been widely applied. Ranft et al. [27] mix over 2 million data in training, LeReS [26] collects over $300$ thousands data, Eftekhar et al. [35] also merge millions of data to build a strong depth prediction model. To train a zero-shot surface normal estimator, Omni-data [35] performs dense reconstruction to generate $14$M frames with surface normal annotations.
@@ 3.1 Ambiguity Issues in Metric Depth Estimation
Figure 5 illustrates an instance of photographs captured using diverse cameras and at varying distances. Solely based on visual inspection, one might erroneously infer that the last two images originate from a comparable location and are captured by the same camera. However, due to differing focal lengths, these images are indeed captured from distinct locations. Consequently, accurate knowledge of camera intrinsic parameters becomes imperative for metric estimation from a single image; otherwise, the problem becomes ill-posed. Recent methodologies, such as MiDaS [27] and LeReS [25], mitigate this metric ambiguity by decoupling metric estimation from direct supervision and instead prioritize learning affine-invariant depth.
@@
Figure 6 (A) depicts the pin-hole perspective projection, where object $A$ located at distance $d_{a}$ is projected to $A^{\prime}$. Adhering to the principle of similarity, it is obvious that:
@@
where $\hat{S}$ and $\hat{S^{\prime}}$ are the real and imaging size respectively. The symbol $\hat{\cdot}$ signifies that variables are expressed in physical metrics (e.g., millimeters). To ascertain $d_{a}$ from a single image, one must have access to the focal length, imaging size of the object, and real-world object size. Estimating the focal length from a single image is challenging and inherently ill-posed. Despite numerous methods having been explored [25, 88], achieving satisfactory accuracy remains elusive. Hereby, we simplify the scenario by assuming known focal lengths for the training/test images. In contrast, understanding the imaging size is much easier for a neural network. To obtain the real-world object size, a neural network needs to understand the semantic scene layout and the object, at which a neural network excels. We define $\alpha=\nicefrac{{\hat{f}}}{{\hat{S^{\prime}}}}$, showing that $d_{a}$ is proportional to $\alpha$.
Figure 6: Pinhole camera model. (A)Object $A$ positioned at a distance $d_{a}$ undergoes projection onto the image plane. (B) Employing two cameras for capturing an image of the car. The left one has a larger pixel size. Although the projected imaging sizes are the same, the pixel-represented images (resolution) are different.
We observe the following regarding sensor size, pixel size, and focal length.
O1: Sensor size and pixel size do not affect metric depth estimation. Based on perspective projection (Fig. 6 (A)), sensor size only influences the field of view (FOV) and is not relevant to $\alpha$, hence it does not affect metric depth estimation. For pixel size, consider two cameras with different pixel sizes ($\delta_{1}=2\delta_{2}$) but the same focal length $\hat{f}$, capturing the same object at distance $d_{a}$. Fig. 6 (B) displays their captured images. According to the preliminaries, the pixel-represented focal length is $f_{1}=\frac{1}{2}f_{2}$. Since the second camera has smaller pixel sizes, the resolution of the pixel-represented image is given by $S^{\prime}_{1}=\frac{1}{2}S^{\prime}_{2}$, despite both having the same projected imaging size $\hat{S^{\prime}}$. According to Eq. (1), we have $\frac{\hat{f}}{\delta_{1}\cdot S^{\prime}_{1}}=\frac{\hat{f}}{\delta_{2}\cdot S^{\prime}_{2}}$, which implies $\alpha_{1}=\alpha_{2}$ and $d_{1}=d_{2}$. This means that variations in camera sensors do not impact the estimation of metric depth.
O2: The focal length is vital for metric depth estimation. Figure 5 shows the challenge of metric ambiguity caused by an unspecified focal length, which is further discussed in Figure 7. In the scenario where two cameras ($\hat{f}_{1}=2\hat{f}_{2}$) are positioned at distances $d_{1}=2d_{2}$, the imaging sizes remain consistent for both cameras. As a result, the neural network struggles to distinguish between different supervision labels based solely on visual cues. To address this issue, we propose a canonical camera transformation method to reduce conflicts between supervision requirements and image representations.
@@
Unlike depth, surface normal does not have any metric ambiguity problem. In Fig. 8, we illustrate this concept with two depth maps at varying scales, denoted as $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, featuring distinct metrics d1 and d2, respectively, where $d_{1}<d_{2}$. After upprojecting the depth to the 3D point cloud, the dolls are in different depths $d_{1}$ and $d_{2}$. However, the surface normals $\mathbf{n}_{1}$ and $\mathbf{n}_{2}$ corresponding to a certain pixel $A^{\prime}\in\mathbf{I}$ remain the same.